OpenStudy (anonymous):

Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 6)(x − 6) and (x − 6)(x − 6)

OpenStudy (anonymous):

(x+6)(x-6)=x^2-36 special factoring, the difference of squares

OpenStudy (anonymous):

(x-6)(x-6) =x^2-12x+36, special factoring, perfect square trinomial

OpenStudy (anonymous):

Compare: They both have perfect squares. They both are special factoring cases.

OpenStudy (anonymous):

Contrast. The difference of squares the middle terms cancel out. The perfect square trinomial the middle terms dont cancel.

OpenStudy (anonymous):

@puertoricangirl96